Can someone help me solve this problem using the Substitution or Elimination Method?
y=5x+2 3x-y=-13

Answer by checkley71(8403)

Y=5X+2 NOW SUBSTITUTE (5X+2) FOR Y IN THE OTHER EQUATION.
3X-(5X+2)=-13
3X-5X-2=-13
-2X=-13+2
-2X=-11
X=-11/-2
X=5.5 ANSWER.
Y=5*5.5+2
Y=27.5+2
Y=29.5 ANSWER.
PROOF
3*5.5-29.5=-13
16.5-29.5=-13
-13=-13